# Zero- to low-field -spectroscopy with a diamond magnetometer

### v2

@@ Abstract
We report measurements of zero- to ultra-low-field nuclear magnetic resonance (ZULF NMR) signals at frequencies of a few hertz with a diamond-based magnetic sensor. The sensing diamond is a truncated pyramid with height and a base. The minimum stand-off distance is mm, and the sensor sensitivity is 13 pT/ at frequencies above 5 Hz with -like behavior at lower frequencies. NMR signals were generated via signal amplification by reversible exchange (SABRE) parahydrogen-based hyperpolarization resulting in zero-field signals at 1.7 Hz and 3.4 Hz corresponding to the expected hetero-nuclear -coupling pattern of acetonitrile. This work demonstrates a magnet-free platform for detecting chemically specific NMR signals at ultra-low frequencies paving the way for portable noninvasive diagnostics in microscopic sample volumes for biomedicine, industrial sensing through metal enclosures, and field-deployable quantum analytical devices.
@@ I Introduction
A sensor to be used for the magnetic detection of ZULF NMR spectra of miniature samples ideally needs to be small (¡1 mm) to enable a short stand-off distance, have high sensitivity (in the pT/ range or better), and a wide bandwidth (on the order of 1 kHz). Currently, sensors for ZULF NMR include superconducting quantum interference devices (SQUIDs) Buckenmaier et al. (2017), atomic magnetometers Ledbetter et al. (2009), and magnetoresistive (MR) sensors Picazo-Frutos,Román et al. (2025). An alternative strategy is shuttling the sample, after evolution under ZULF conditions, into a high magnetic field, allowing for detection by inductive coils Barskiy et al. (2025). However, the need for cryogenic cooling limits the usability of SQUIDs. Magnetoresistive sensors, while potentially enabling higher spatial resolution, present challenges due to their residual magnetic fields, which can interfere with the ZULF conditions required for spectral acquisition Picazo-Frutos,Román et al. (2025). Optically pumped atomic magnetometers (OPM) can, in principle, achieve high spatial resolution on the order of hundreds of micrometers Fang,Xiujie et al. (2020), along with sensing bandwidths extending from DC to several hundred kilohertz Li et al. (2020). However, the sensitivity is rapidly lost for small sensor volumes, and the stand-off distance is limited by the vapor cell itself.
@@ II NV center zero bias field magnetometry
One of the most common protocols in NV magnetometry is optically detected magnetic resonance (ODMR) Barry et al. (2020). In this method, the diamond electron spins are polarized by optical pumping into the state in the ground-state electron spin-1 manifold. A microwave (MW) magnetic field is applied to drive transitions within the NV ground-state manifold. These transitions move the NV spins into one or both of the depopulated sublevels, which experience Zeeman frequency shifts under a magnetic field. Transitions can be optically detected Barry et al. (2020) as the photoluminescence is lower for these states compared to . For a single NV center photoluminescence reduces by 30 % Barry et al. (2020).
@@
Several techniques have been developed to enable zero-field sensing with NV centers. One approach uses circularly polarized MW fields Zheng et al. (2020); Lenz et al. (2021) to selectively excite one of the degenerate sublevels. However, this method is technically challenging because of the difficulty of generating circular polarization along all four NV axes and achieving uniform circularity across the entire diamond volume. Another technique leverages cross-relaxation features near zero magnetic field, which are prominent in NV centers without preferential orientation Dhungel,Omkar et al. (2024). However, this method has not yet reached sensitivities comparable to standard finite-field ODMR. Magnetically sensitive noise floors have been demonstrated on the order of around 100 nT/ with 2 nT/ photon shot noise Dhungel et al. (2024). The method still requires the application of a DC or modulated bias field. Hyperfine coupling with 13C nuclei has also been used for zero-field sensing in single NV centers Wang,Ning et al. (2022). However, extending this to NV ensembles requires the use of isotopically enriched 13C diamond, which increases electron spin dephasing and degrades sensitivity. Furthermore, ensemble averaging over varying NV–13C distances introduces inhomogeneous effective fields that complicate interpretation.
In this study, we employ a zero-field magnetometry protocol with a modulated bias (6.12 kHz along [100] crystallographic direction with T amplitude) achieving sensitivities of 13 pT/ comparable to magnetically-biased ODMR Sekiguchi,N. et al. (2024); Zheng et al. (2020). The presence of a T-scale magnetic field broadens the zero-field ODMR feature, which, in turn, reduces contrast. This change in peak amplitude can itself be viewed as a form of magnetic resonance, with a peak at zero bias field utilizing all NV axes. Therefore, by modulating and demodulating the magnetic field, a magnetic dispersive signal can be extracted. This signal features a zero crossing at zero bias, allowing for high-sensitivity sensing with a linear response in this regime. This method was already demonstrated and described analytically as an alternative to zero-field sensing using circularly polarized microwaves Zheng et al. (2020).
While the zero-field feature frequency is independent to magnetic field changes around zero bias-magnetic fields, temperature variations still result in frequency shifts Shim et al. . To avoid systematic effects due to these shifts, we stabilize the MW to the zero-field feature. For this we generate an error signal by sinusoidally frequency modulating the MW frequency (modulation frequency 140 kHz, modulation amplitude 100 kHz) and demodulating it with a lock-in amplifier. The demodulated signal is linear to frequency shifts due to temperature variation and just quadratic to frequency shifts due to magnetic field changes. Stabilizing the MW frequency via a proportional-integral-differential (PID) control loop makes the system largely insensitive to temperature fluctuations, enabling long-term measurement campaigns.
@@ III.1 Hyperpolarization
ZULF NMR signals originate from [15N]-acetonitrile molecules. A pulse of 860 s with 6 T amplitude is applied to initiate the free magnetization decay. The diamond sensor data acquisition lasted 40 s, initiated by a software trigger synchronized with the ZULF sequence. All chemicals except for the catalyst were acquired through Sigma-Aldrich. The catalyst was prepared based on previously developed methods Vazquez-Serrano et al. (2006); Savka and Plenio (2015); Blanchard et al. (2021b).
III.2 Diamond sensor
@@
The principle of the sensor is based on the ODMR technique Barry et al. (2020). The sensor head is designed to optimize light collection Omar,Muhib et al. (2023), see Fig. 2. It includes an output coupler for the optical fiber delivering the pump laser light, and producing a beam of 300 m in diameter. This beam passes through a polarizing beam splitter to clean its polarization. A beam sampler captures part of the light before the beam propagates through the backside of a dichroic mirror and into a lens that focuses the light into the sensing diamond (containing 3.7 ppm of NV and 10 ppm of nitrogen). The sensing diamond is shaped like a truncated pyramid for side collection and mounted on a diamond anvil to enhance the collection efficiency by a factor of four Omar,Muhib et al. (2023). In addition, a silver reflective coating was applied on the diamond components to avoid leakage of the green laser light and to increase light collection efficiency further. The red photoluminescence (PL) emitted by the diamond is collected with the same lens that focuses the laser beam. This light is reflected by the front side of the dichroic mirror and directed to a photodiode (PD). The sensor uses non-magnetic Hamamatsu photodiodes (S13228-01) soldered onto a printed circuit board (PCB). The current from the green pickup beam PD and the red light PD are subtracted directly on the PCB for differential detection to reduce the sensitivity to laser intensity variation. A 3D-printed clipping mask is used to partially cover the green pick up beam PD , allowing fine adjustment of the differential detection scheme Sekiguchi,N. et al. (2024). With 150 mW of green laser light power focused on the diamond, 5 mW of red photoluminescent light is collected on the photodiode. The mobile sensor setup includes an instrument rack on wheels. This rack includes a multi-frequency lock-in amplifier (LIA), a breadboard with a Verdi G5 laser (532 nm) coupled into a high-power single-mode optical fiber that delivers light to the sensor head, a preamplifier to enhance the signal from the sensor head, and a microwave signal generator. The output of the microwave generator is mixed with that of an RF source and then amplified by a microwave amplifier. This frequency-mixing technique alters the hyperfine-resolved spectrum by simultaneously driving two resonance transitions with two microwave frequencies and therefore increases the optical contrast of the electron spin resonance. This in turn directly enhances the magnetic measurement sensitivity Barry et al. (2020). The microwaves are delivered to the sensor via a flexible PCB. The differential signal of the sensor head is sent to the preamplifier and subsequently to the lock-in amplifier. Around the diamond, a coil is mounted, and its current is modulated with one of the lock-in amplifier outputs to enable the magnetic sensing protocol. The custom designed PCBs for the photodiode and microwave delivery are discussed in the Appendix.
@@
III.3 The vapor cell magnetometer
@@ VI Acknowledgments
This work was supported by the EU, project HEU-RIA-MUQUABIS-101070546, by the Helmholtz Association project Quantum Sensing for Fundamental Physics (QS4Physics) from the Innovation pool of the research field Helmholtz Matter, by the DFG, project FKZ: SFB 1552/1 465145163, by the Alexander von Humboldt Foundation in the framework of the Sofja Kovalevskaja award and by the German Federal Ministry of Research, Technology and Space (BMFTR) within the Quantumtechnologien program via the DIAQNOS project (project no. 13N16455). We thank Hamamatsu for providing the photodiodes as samples.
@@ VII.1 Diamond sensor design choices
The polycrystalline diamond plate was custom-fabricated by Medidia Tec GmbH and assembled with the other diamond components, the science diamond [Element Six (UK) Ltd.] and the diamond anvil (Dutch Diamond Technologies BV) at their facility.

### v3

# Zero- to ultralow-field -spectroscopy with a diamond magnetometer

@@ Abstract
We report measurements of zero- to ultralow-field nuclear magnetic resonance (ZULF NMR) signals at frequencies of a few hertz with a diamond-based magnetic sensor. The sensing diamond is a truncated pyramid with height and a base. The minimum stand-off distance is mm, and the sensor sensitivity is 13 pT/ at frequencies above 5 Hz with -like behavior at lower frequencies. NMR signals were generated via signal amplification by reversible exchange (SABRE) parahydrogen-based hyperpolarization resulting in zero-field signals at 1.7 Hz and 3.4 Hz corresponding to the expected hetero-nuclear -coupling pattern of acetonitrile. This work demonstrates a magnet-free platform for detecting chemically specific NMR signals at ultra-low frequencies paving the way for portable noninvasive diagnostics in microscopic sample volumes for biomedicine, industrial sensing through metal enclosures, and field-deployable quantum analytical devices.
@@ I Introduction
A sensor to be used for the magnetic detection of ZULF NMR spectra of miniature samples ideally needs to be small (1 mm) to enable a short stand-off distance, have high sensitivity (in the pT/ range or better), and a wide bandwidth (on the order of 1 kHz). Currently, sensors for ZULF NMR include superconducting quantum interference devices (SQUIDs) Buckenmaier et al. (2017), atomic magnetometers Ledbetter et al. (2009), and magnetoresistive (MR) sensors Picazo-Frutos,Román et al. (2025). An alternative strategy is shuttling the sample, after evolution under ZULF conditions, into a high magnetic field, allowing for detection by inductive coils Barskiy et al. (2025). However, the need for cryogenic cooling limits the usability of SQUIDs. Magnetoresistive sensors, while potentially enabling higher spatial resolution, present challenges due to their residual magnetic fields, which can interfere with the ZULF conditions required for spectral acquisition Picazo-Frutos,Román et al. (2025). Optically pumped atomic magnetometers (OPM) can, in principle, achieve high spatial resolution on the order of hundreds of micrometers Fang,Xiujie et al. (2020), along with sensing bandwidths extending from DC to several hundred kilohertz Li et al. (2020). However, the sensitivity is rapidly lost for small sensor volumes, and the stand-off distance is limited by the vapor cell itself.
@@ II NV center zero bias field magnetometry
One of the most common protocols in NV magnetometry is optically detected magnetic resonance (ODMR) Barry et al. (2020). In this method, the diamond electron spins are polarized by optical pumping into the state in the ground-state electron spin-1 manifold. A microwave (MW) magnetic field is applied to drive transitions within the NV ground-state manifold. These transitions move the NV spins into one or both of the depopulated sublevels, which experience Zeeman frequency shifts under a magnetic field. Transitions can be optically detected Barry et al. (2020) as the photoluminescence (PL) is lower for these states compared to . For a single NV center PL reduces by 30 % Barry et al. (2020).
@@
Several techniques have been developed to enable zero-field sensing with NV centers. One approach uses circularly polarized MW fields Zheng et al. (2020); Lenz et al. (2021) to selectively excite one of the degenerate sublevels. However, this method is technically challenging because of the difficulty of generating circular polarization along all four NV axes and achieving uniform circularity across the entire diamond volume. Another technique leverages cross-relaxation features near zero magnetic field, which are prominent in NV centers without preferential orientation Dhungel,Omkar et al. (2024).
However, this method has not yet reached sensitivities comparable to standard finite-field ODMR. Magnetically sensitive noise floors have been demonstrated on the order of around 100 nT/ with 2 nT/ photon shot noise Dhungel et al. (2024). The method still requires the application of a DC or modulated bias field. Hyperfine coupling with 13C nuclei has also been used for zero-field sensing in single NV centers Wang,Ning et al. (2022). However, extending this to NV ensembles requires the use of isotopically enriched 13C diamond, which increases electron spin dephasing and degrades sensitivity. Furthermore, ensemble averaging over varying NV–13C distances introduces inhomogeneous effective fields that complicate interpretation.
In this study, we employ a zero-field magnetometry protocol with a modulated bias (6.12 kHz along [100] crystallographic direction with T amplitude) achieving sensitivities of 13 pT/ comparable to magnetically-biased ODMR Sekiguchi,N. et al. (2024); Zheng et al. (2020). This approach provides high-sensitivity sensing without requiring circularly polarized microwaves, as demonstrated and analyzed previously Zheng et al. (2020). While the zero-field feature frequency is independent to magnetic field changes around zero bias-magnetic fields, temperature variations still result in frequency shifts Shim et al. . To avoid systematic effects due to these shifts, we stabilize the MW to the zero-field feature. Further details can be found in the Method section.
@@ III.1 Hyperpolarization
ZULF NMR signals originate from [15N]-acetonitrile molecules. In the absence of a bias field, the main frequency components of the signal are determined by the -coupling constant of Hz (Fig. 1a), resulting in two principal transitions within the zero-bias energy-level structure (see Fig. 1b) of the compound: one at the -frequency and the other at twice that frequency. These transitions can be magnetically measured and are visible in the amplitude spectrum of the magnetization decay of the sample (Fig. 1c).
@@
A pulse of 860 s with 6 T amplitude is applied to initiate the free magnetization decay. The diamond sensor data acquisition lasted 40 s, initiated by a software trigger synchronized with the ZULF sequence. All chemicals except for the catalyst were acquired through Sigma-Aldrich. The catalyst was prepared based on previously developed methods Vazquez-Serrano et al. (2006); Savka and Plenio (2015); Blanchard et al. (2021b).
III.2 Zero-field sensing and microwave locking
In the following we discuss the microwave locking technique and the zero-field sensing methodology. To understand the mechanism it is sufficient to consider the ODMR signal for a single axis neglecting hyperfine structure under MW driving at frequency . With a subtracted background PL the signal can be modeled proportional to the sum of two Lorentzian functions. One for each ground state transition :
where is the full-width-at-half-maximum (FWHM), is the linearly approximated temperature dependent zero-field ground state transition frequency with temperature independent offset frequency, MHz and proportionality constant . In general is temperature dependentRodzoń et al. (2025), but at room temperature, Acosta et al. (2010). The Larmor frequency is a linear shift of the resonance frequency as a function of the applied on-axis magnetic field which changes sign depending on the level. The NV center’s gyromagnetic ratio is GHz/T.
At T the two transitions become degenerate. Phase-sensitive detection of with a lock-in amplifier (LIA) for a small modulation of the MW frequency results in the derivative signal of Eq. (1) with respect to the microwave frequency . Without background magnetic field this signal is dispersive around and can be used to lock the microwave frequency.
The derivative can be evaluated:
where is a proportionality constant depending on the lock-in settings, modulation parameters and resonance properties. The Zeeman shift cancels itself and linear sensitivity to temperature is retained. Similarly, for a small modulation of the magnetic field around zero field, the derivative with respect to the offset magnetic field can be computed using phase-sensitive detection. To first order this is linearly dependent on the background magnetic field :
where is a proportionality constant depending on modulation parameter, lock-in settings and resonance properties. By simultaneously modulating both MW frequency and magnetic field and demodulating at separate reference frequencies, two independent dispersive signals can be obtained: one sensitive primarily to MW-frequency shifts (e.g., due to temperature drifts) and one linear in magnetic field. Care has to be taken, that the frequencies are sufficiently different, such that there are no cross-modulation features in the frequency range of the magnetometer.
This forms the basis for high-sensitivity zero-field magnetometry using NV centers. In practice, we generate an error signal for the microwave locking by sinusoidally modulating the MW frequency (modulation frequency 140 kHz, amplitude 100 kHz) and demodulating it with a lock-in amplifier. The magnetically sensitive signal similarly results from sinusoidally modulating the magnetic field (modulation frequency 6.12 kHz, amplitude T) and demodulating it with a second lock-in amplifier. Stabilizing the MW frequency using a proportional–integral–derivative (PID) control loop thus makes the system largely insensitive to temperature fluctuations, enabling long-term measurements, while the output of the second lock-in amplifier is a direct measure of the magnetic field at the position of the diamond.
III.3 Diamond sensor
@@
The principle of the sensor is based on the ODMR technique Barry et al. (2020). The sensor head is designed to optimize light collection Omar,Muhib et al. (2023), see Fig. 2. It includes an output coupler for the optical fiber delivering the pump laser light, and producing a beam of 300 m in diameter. This beam passes through a polarizing beam splitter to clean its polarization. A beam sampler captures part of the light before the beam propagates through the backside of a dichroic mirror and into a lens that focuses the light into the sensing diamond (containing 3.7 ppm of NV and 10 ppm of nitrogen). The sensing diamond is shaped like a truncated pyramid for side collection and mounted on a diamond anvil to enhance the collection efficiency by a factor of four Omar,Muhib et al. (2023). In addition, a silver reflective coating was applied on the diamond components to avoid leakage of the green laser light and to increase light collection efficiency further. The red PL emitted by the diamond is collected with the same lens that focuses the laser beam. This light is reflected by the front side of the dichroic mirror and directed to a photodiode (PD). The sensor uses non-magnetic Hamamatsu photodiodes (S13228-01) soldered onto a printed circuit board (PCB). The current from the green pickup beam PD and the red light PD are subtracted directly on the PCB for differential detection to reduce the sensitivity to laser intensity variation. A 3D-printed clipping mask is used to partially cover the green pick up beam PD , allowing fine adjustment of the differential detection scheme Sekiguchi,N. et al. (2024). With 150 mW of green laser light power focused on the diamond, 5 mW of red photoluminescent light is collected on the photodiode. The mobile sensor setup includes an instrument rack on wheels. This rack includes a multi-frequency lock-in amplifier (Zurich Instruments AG, HF2LI), a breadboard with a Coherent, Verdi G5 laser (532 nm) coupled into a high-power single-mode optical fiber that delivers light to the sensor head, a preamplifier to enhance the signal from the sensor head, and a microwave signal generator. The output of the microwave generator Rohde Schwarz, SMA100b is mixed with that of an RF source (Keysight Technologies, 33512B) and then amplified by a microwave amplifier (Mini-circuits, ZHL-16W-43-S+). This frequency-mixing technique alters the hyperfine-resolved spectrum by simultaneously driving two resonance transitions with two microwave frequencies and therefore increases the optical contrast of the electron spin resonance. This in turn directly enhances the magnetic measurement sensitivity Barry et al. (2020). The microwaves are delivered to the sensor via a flexible PCB. The differential signal of the sensor head is sent to the preamplifier and subsequently to the lock-in amplifier. Around the diamond, a coil is mounted, and its current is modulated with one of the lock-in amplifier outputs to enable the magnetic sensing protocol. The custom designed PCBs for the photodiode and microwave delivery are discussed in the Appendix.
@@
III.4 The vapor cell magnetometer
@@ VI Acknowledgments
This work was supported by the EU, project HEU-RIA-MUQUABIS-101070546, the Initiative and Networking Fund of the Helmholtz Association (Project No. VH-NG-20-20) by the Helmholtz Association project Quantum Sensing for Fundamental Physics (QS4Physics) from the Innovation pool of the research field Helmholtz Matter, by the DFG, project FKZ: SFB 1552/1 465145163, by the Alexander von Humboldt Foundation in the framework of the Sofja Kovalevskaja award and by the German Federal Ministry of Research, Technology and Space (BMFTR) within the Quantumtechnologien program via the DIAQNOS project (project no. 13N16455). We thank Hamamatsu for providing the photodiodes as samples.
@@ VII.1 Diamond sensor design choices
The polycrystalline diamond plate was custom-fabricated by Medidia Tec GmbH and assembled with the other diamond components, the science diamond [Element Six (UK) Ltd.] and the diamond anvil (Dutch Diamond Technologies BV) at their facility. The sensing diamond was fixed to the diamond anvil using Norland Optical Adhesive 170.